The age of a parent and child add up to 39 years,In 3years the parent will be 4 times as old as the child.find the age of the parent and the child.
Answer by ankor@dixie-net.com(22740)

let p = age of parent
let c = age of child
:
Write an equation for each statement
:
" The age of a parent and child add up to 39 years,"
p + c = 39:
or
p = (39-c); use this form for substitution
;
"In 3years the parent will be 4 times as old as the child."
add 3 years to each
p + 3 = 4(c+3)
p + 3 = 4c + 12
p = 4c + 12 - 3
p = 4c + 9
Replace p with (39-c) from the 1st statement
39 - c = 4c + 9
39 - 9 = 4x + c
30 = 5c
c = 30/5
c = 6 yrs old is the child
:
I'll let you find the parent's age, check your solutions in the 2nd statement